The length of a rectangle is 4 inches more than three times the width. The perimeter is 88 inches. Find the length and width.
Answer by Maths68(1474)

Perimeter = p = 88 inches
Width = w = ?
Length = l = 3w+4

Perimeter of rectangle = 2 (l+w)
Put l = 3w+4
p=2(3w+4+w)
88=2(4w+4)
88/2=2(4w+4)/2
44=4w+4
44=4(w+1)
44/4=4(w+1)/4
11=w+1
11-1=w
10=w

Perimeter = p = 88 inches
Width = w = 10 inches
Length = l = 3w+4 = (3)(10)+1 = 30+4 = 34 inches